The drug Prevnar is a vaccine meant to prevent certain types of bacterial meningitis. It is typically administered to infants starting around two months old. In a randomized double-blind clinical trials of Prevnar, infants were randomly divided into two groups. Group 1 received Prevnar while Group 2 received a control vaccine. After the second dose, 137 of 452 subjects in group 1 effect experienced drowsiness as a side effect. After the second dose, 31 of 9 in Grou 4. 9 subjects p 2 experienced drowsiness as a side effect. Does the evidence suggest that there is a difference proportion between the groups at a 0.05 level of significance.
Answer #1

H0: P1 = P2

H1: P1 eq P2

widehat p1 = 137/452 = 0.303

02 = 31/99 = 0.313

The pooled proportion(P) = (widehat p1 * n1 + 02 * n2)/(n1 + n2)

                                         = (0.303 * 452 + 0.313 * 99)/(452 + 99)

                                         = 0.305

SE = sqrt(P(1 - P)(1/n1 + 1/n2))

     = sqrt(0.305 * (1 - 0.305) * (1/452 + 1/99))

    = 0.051

The test statistic z = (widehat p1 - widehat p2)/SE

                             = (0.303 - 0.313)/0.051

                             = -0.20

P-value = 2 * P(Z < -0.20)

             = 2 * 0.4207

             = 0.8414

Since the P-value > alpha(0.8414 > 0.05), so we should not reject the null hypothesis.

So at alpha = 0.05, there is not sufficient evidence to conclude that there is a difference proportion between the groups.